# Numerical Aspects of Large Deviations

### v1

@@ 5 Conclusion
This text provides the fundamentals to access the tails of distributions in simulations of stochastic systems. Then main idea of the central approach is to separate the randomness from the othergredients of the model and keep the randomness in a vector of $U(0,1)$ distributed entries. This allows one to perform a Markov Chain Monte Carlo simulation including a bias, which drives the simulation to the desired range of values. Particularily useful is the exponential bias, with a temperature-like parameter controlling the range of values. By performing simulation for different temperatures, often the distribution can be obtained numerically over hundres of decades in probability. Here, the methods are illustrated using a very simple model, the Bernoulli process. Since any stochastic simulation relies in principle on $U(0,1)$ random numbers, the approach is very general and can be applied to a wide range of equilibrium and non-equilibrium models. Still, many issues can arise, like chosing a suitable bias or to determine efficient moves when cahnhing the configurations in the Markov chain.
@@ 6 Acknowledgements
The authors thanks Abishek Dhar, Joachim Krug, Satya N. Majumdar, Alberto Rosso, and Grégory Schehr for organizing the summer school “Theory of Large Deviations and Applications” in Les Houches 2024. The author is thankful to Peter Werner for critically reading the manuscript. He furthermore is grateful to Winfried G. Schneeweiß for introducing him to the method of biased simulations which happened during the author’s Diploma thesis in Computer Science at the Fernuniversität Hagen (Germany) in 1992. He also thanks Peter Grassberger for introducing him to the field of the statistics of biological sequence alignment during a workshop at Santa Barbara (USA) in 2001, which motivated the first physics large-deviation project [34] of the author. The author finally is much obliged to all his collaborators in the field, which have become too many to list them all here.

### v2

@@ 4.4.6 Direct encoding of configurations
4.5 Computational resources
One might wonder what computational resources are needed to study the tails of the distribution. First, most of the time is spent in running single instances of the model of interest, for a given set of random numbers, or for a given directly encoded instance. This sets the time scale needed to perform the MCMC simulations, say for an exponential bias $\exp(-S/\Theta)$ with a given temperature $\Theta$. A typical number of MC steps is $O(10^{6})$. Often a run for a single instance requires only few milliseconds, if order of 1000 variables describe the instance. This means the biased MCMC will run few 1000s, which can be done on a laptop, as it is the case for the simple Bernoulli process considered here. If one single runs requires one second, the MCMC will take correspondingly more.
Second, one has to take into account how many of temperature values $\Theta$ have to be considered, since the simulations have to be performed independently for all of them. Typically, to reach probabilities as small as $10^{-50}$, of the order of 10 temperatures are needed. If one wants to go much lower, many more temperatures are needed. For example in the study of the directed polymers in disordered media [14, 15, 46], probabilities as small as $10^{-1000}$ were obtained, which required more than 200 different values of $\Theta$. Since also the system sizes were rather large, leading to longer running times to evaluate single instances, a parallel implementation on a high-performance cluster was used and required several weeks while running on few hundreds of cores in parallel.
4.6 Other approaches
The presented approach of changing random numbers and using an exponential bias is rather general but not completely universal, so other approaches exist.
Sometimes, in particular if $P(S)$ is not concave, the exponential bias $\exp(-S/\Theta)$ leads to an effective distribution $P(S)\exp(-S/\Theta)/Z(\Theta)$ which is not bell-like around a typical value as shown in Fig. 7, but instead could exhibit a two-peak structure [39]. In this case the values of $S$ between the two peaks are hard to access. This might turn even impossible if the system size is large. One way out could be to use a more concentrated bias, which forces the biased distribution to be bell-like. One could, e.g., use a Gaussian bias [48] $\exp(-(S-S_{0})^{2}/(2\sigma^{2}))$, as it has recently been applied to study the distribution of susceptibilities for diluted magnetic systems [49].
An even more general approach is to aim at using $1/P(S)$ as the bias. Since the resulting distribution is proprtional to $P(S)/P(S)=1$, i.e., one obtains a more or less uniform sampling. This approach is called Umbrella sampling [50]. Nevertheless, typically $P(S)$ is now known. In this case one tries to construct $P(S)$ iteratively with increasing accurcy during the simulation, starting from a flat distribution. For this purposes methods like the Wang-Landau approach are useful [51], which was applied to study rare trajectories of infection dynamics [29, 30, 31, 32], as already mentioned above. A related algorithm is the Multi-canonical Ensemble [52] which has been applied for the rare-event case to testing error-correcting codes [53]. Another issue could be that the system of interest is simulated with a given package, which is hard or even impossible to change, such that one cannot replace the internal calls to a random number generator by accesses to a vector of random numbers. Also, sometimes the actual dynamics is free of randomness, i.e., deterministic. In both cases, one could perform small random changes to the initial or intermediate states of the system under consideration. Such approaches are used for rare-event sampling of climate models [54]. Here, the single runs of the original model are so demanding that indeed a MCMC sampling with each step being a full run of the climate model over the full period of interest would require too much time. Here, in particular in the case where the quantity $S$ of interest is integrated over time., e.g., the physical temperature in the climate model, population based approaches, so called Cloning Algorithms [55], sometimes referred as genetic selection algorithms, are convenient. Here, instances with slightly different initial conditions are simulated, and after a suitable number of steps for the model, the population is reshaped according the biased probabilities of the different population members. Such approaches have also been used for some models, like the Totally Asymmetric Exclusion Process (TASEP) [56].
For other problems, sometime one is not interested in particular rare endpoints of a process, but in rare trajectories between known endpoints. This is the case when considering configurational changes of proteins. Here the Transition-Path Sampling Approach [57, 58] is very useful.
@@ 5 Conclusion
This text provides the fundamentals to access the tails of distributions in simulations of stochastic systems. Then main idea of the central approach is to separate the randomness from the othergredients of the model and keep the randomness in a vector of $U(0,1)$ distributed entries. This allows one to perform a Markov Chain Monte Carlo simulation including a bias, which drives the simulation to the desired range of values. Particularily useful is the exponential bias, with a temperature-like parameter controlling the range of values. By performing simulation for different temperatures, often the distribution can be obtained numerically over hundreds of decades in probability. Here, the methods are illustrated using a very simple model, the Bernoulli process. Since any stochastic simulation relies in principle on $U(0,1)$ random numbers, the approach is very general and can be applied to a wide range of equilibrium and non-equilibrium models. Still, many issues can arise, like choosing a suitable bias or to determine efficient moves when changing the configurations in the Markov chain.
@@ 6 Acknowledgements
The authors thanks Abishek Dhar, Joachim Krug, Satya N. Majumdar, Alberto Rosso, and Grégory Schehr for organizing the summer school “Theory of Large Deviations and Applications” in Les Houches 2024 and for inviting me to give a series of lectures which was the basis of the current texte. The author is thankful to Peter Werner for critically reading the manuscript. He furthermore is grateful to Winfried G. Schneeweiß for introducing him to the method of biased simulations which happened during the author’s Diploma thesis in Computer Science at the Fernuniversität Hagen (Germany) in 1992. He also thanks Peter Grassberger for introducing him to the field of the statistics of biological sequence alignment during a workshop at Santa Barbara (USA) in 2001, which motivated the first physics large-deviation project [37] of the author. The author finally is much obliged to all his collaborators in the field, which have become too many to list them all here.